### v2

# Fault-tolerant fermionic quantum computing

**footnotetext: Correspondence to: schuckertalexander@gmail.com, mgullans@umd.edu
Fault-tolerant fermionic quantum computing
[…]
(July 16, 2025)
Simulating the dynamics of electrons and other fermionic particles in quantum chemistry [1], materials science [2], and high-energy physics [3] is one of the most promising applications of fault-tolerant quantum computers. However, the overhead in mapping time evolution under fermionic Hamiltonians to qubit gates [4, 5, 6, 7] renders this endeavor challenging [8, 9]. We introduce fermion-qubit fault-tolerant quantum computing, a framework which removes this overhead altogether. Using native fermionic operations we first construct a repetition code which corrects phase errors only. Within a fermionic color code, which corrects for both phase and loss errors, we then realize a universal fermionic gate set, including transversal fermionic Clifford gates. Interfacing with qubit color codes we introduce qubit-fermion fault-tolerant computation, which allows for qubit-controlled fermionic time evolution, a crucial subroutine in state-of-the-art quantum algorithms for simulating fermions [10, 11, 12, 13]. As an application, we consider simulating crystalline materials [12], finding an exponential improvement in circuit depth for a single time step from $\mathcal{O}(N)$ to $\mathcal{O}(\log(N))$ with respect to lattice site number $N$ while retaining a fermion-site count of $\tilde{\mathcal{O}}(N)$, implying a linear-in-$N$ end-to-end gate depth for simulating materials, as opposed to quadratic in previous approaches. We also introduce a fermion-inspired qubit algorithm with $O(\mathrm{log}(N)$ depth, but which uses a prohibitive number of additional ancilla qubits compared to fermionic hardware. We show how our framework can be implemented in neutral atoms, overcoming the apparent inability of neutral atoms to implement non-number-conserving gates. Our work opens the door to fermion-qubit fault-tolerant quantum computation in platforms with native fermions such as neutral atoms, quantum dots and donors in silicon, with applications in quantum chemistry, material science, and high-energy physics.
Figure 1: Logical fermions and architecture. a, A logical qubit is encoded in a codeblock formed by many physical qubits. Similarly, we encode a logical fermion into a codeblock formed by many physical fermions. While logical qubit operators commute, logical fermionic operators anti-commute between two codeblocks. Physical fermions can be encoded in fermionic neutral atoms trapped in optical tweezers or lattices (indicated by hour-glass shapes) or in electrons trapped by electric fields in quantum dots or donors in silicon (Methods). b, Neutral-atom architecture, circuit notation, and universal qubit-fermion gate set. Single-fermion gates are performed by detuning the tweezer potential; single-qubit gates by Raman lasers; $\mathrm{CZ}$ gates by Rydberg lasers; braiding gates by photodissociation of molecules as well as inter-tweezer tunneling. In the braiding gate, the site $i$, c.f. definition in Eq. (4), is indicated by a cross. We distinguish fermion sites from qubits by using dashed lines and drop the f superscript in circuits.
Solving the dynamics of many strongly-interacting fermionic quantum particles is central to understanding the dynamics of quarks in the early universe, the reaction dynamics of molecules, and high-temperature superconductivity. While qubit quantum computers promise to help solve these open problems, simulating the evolution of fermions encoded in qubits encounters a large overhead in both the number of two-qubit gates and the circuit depth [5, 14], which means that a solution of fermionic problems with near-term [8, 9] and even fault-tolerant quantum computers [15, 12] is extremely challenging. Quantum simulation experiments with cold fermionic atoms have made major progress in understanding continuum [16, 17, 18, 19, 20, 21] and lattice systems [22, 23, 24, 25]. These experiments have made steps towards digital computation by trapping individual fermionic atoms in optical tweezers [26, 27]. Additionally, qubits have been encoded into pairs of fermions [28]. Quantum simulations using trapped electrons in the solid state have also been demonstrated [29, 30, 31]. However, turning these fermionic platforms from a tool of scientific exploration into a computational tool requires error-correction and fault-tolerance techniques, for which no practical scheme has been found so far.
Here, we close this gap by introducing a framework for fermion-to-fermion fault-tolerant quantum computation. We first prove a theorem that fermionic error correction is not possible using only number-conserving gates, which are most easily implemented in experimental platforms. We then show how to avoid this theorem by using a non-number-conserving gate and concepts from Majorana error correction. This enables us to explicitly show how to perform a universal fermionic gate set including fault-tolerant transversal gates. To enable usage of state-of-the-art quantum algorithms requiring qubit-controlled gates, we introduce qubit-fermion fault-tolerant computation. We apply our framework to time evolution under the Hamiltonian of a crystalline material and study the advantages of our framework compared to all-qubit quantum computation. Finally, we show how to realize these codes and their operations in neutral atoms by removing the previously assumed restriction to number-conserving gates in neutral atoms and introducing the physics behind the implementation of a non-number-conserving braiding gate.
Fermion-to-fermion error correction gate set
Our central goal is the construction of an error correcting code that encodes a logical fermion, which we define as a creation operator $c^{\dagger}_{a}$ anti-commuting between different codeblocks $a$ and $b$, i.e. $\{c_{a}^{\dagger},c_{b}\}=\delta_{ab}$, see Fig. 1a. This is by contrast to fermionic operators encoded into a single codeblock of a topological error correcting code [32, 7, 33], which are difficult to address via transversal gates [34] and whose efficiency degrades when applied to fermion models that are not local in 2D. Constructing logical operators that anticommute between different codeblocks is impossible using physical qubits because no qubit operator can be constructed which anticommutes between separate codeblocks. Instead, we use physical fermionic degrees of freedom $p^{\dagger}_{i}$ acting on physical sites $i$ and fulfilling $\{p_{i}^{\dagger},p_{j}\}=\delta_{ij}$ to encode logical fermions. A physical fermion can be realized for example by a fermionic atom trapped with light or an electron in a quantum dot. For concreteness, we will focus on the former implementation and comment on a quantum dot implementation in the Methods. As a second crucial ingredient, we interface these fermions with qubits, which are encoded in two internal energy levels of an atom.
We first consider a fermion-number-conserving gate set acting on many fermionic and qubit sites labeled by indices $i,j$,
$\mathrm{T}^{\mathrm{f}}_{i}$ $=\exp\left(i(\pi/4)n_{i}\right),$ (1)
$\mathrm{CZ}^{\mathrm{f}}_{ij}$ $=\exp\left(i\pi n_{i}n_{j}\right),$ (2)
$\mathrm{CZ}^{\mathrm{qf}}_{ij}$ $=\exp\left(i\pi n^{\mathrm{q}}_{i}n_{j}\right),$ (3)
along with the qubit Hadamard gate $\mathrm{H}$, and the qubit phase gate $\mathrm{S}$. For this gate set to be universal (see Methods for a proof), we also need a gate which moves fermions between sites, $\sqrt{\mathrm{iSWAP}}^{\mathrm{f}}_{ij}=\exp\left(i(\pi/4)\left(p_{i}^{\dagger}p_{j}+h.c.\right)\right)$. For convenience, we also define the gates $\mathrm{S}^{\mathrm{f}}_{i}=(\mathrm{T}^{\mathrm{f}}_{i})^{2}$ and $\mathrm{Z}^{\mathrm{f}}_{i}=(\mathrm{S}^{\mathrm{f}}_{i})^{2}$ not necessary for universality. $\mathrm{CZ}^{\mathrm{qf}}$ is a qubit-fermion gate dependent on the fermionic number operator $n_{j}=p_{j}^{\dagger}p_{j}$ and its qubit analogue $n^{\mathrm{q}}_{i}=\ket{1}\bra{1}_{i}$, i.e. the projector on the |$\ket{1}$ state of the qubit. Note that, in principle, we do not require qubits for fermion universality [35], but as we will show, being able to couple fermions to qubits will lead to key advantages. Neutral-atom platforms are considered to be constrained to this number-conserving gate set [36] as usually the atom number is conserved during gates and each tweezer is initially prepared in a number eigenstate.
The first question is therefore whether these number-conserving operations suffice to encode logical fermions, which is equivalent to demanding that the codestates are eigenstates of the total fermion number operator. Unfortunately, such states are not sufficient:
Theorem.
An error-correcting code using eigenstates of the total fermion number operator as codestates does not have logical fermion operators.
The proof (see Methods) relies on the fact that codestates that are eigenstates of the total fermion number operator need to be eigenstates with the same number eigenvalue due to the Knill-Laflamme conditions. This means that logical operators cannot change the total fermion number and therefore need to be even weight in fermionic creation/annihilation operators. Hence, logical operators are bosonic.
To circumvent this theorem, we use fermion parity eigenstates as codestates, i.e. states that are superpositions of states that contain either only an odd or only an even number of fermions. A universal gate set on such states requires a non-number-conserving gate. In analogy to the braiding operation performed in Majorana nanowires [37], we define the braiding gate
$\mathrm{BRAID}_{ij}=\exp\left(i\frac{\pi}{4}\left(p_{i}^{\dagger}-p_{i}\right)\left(p_{j}^{\dagger}+p_{j}\right)\right),$ (4)
which will also turn out to be a natural logical gate in our codes. We propose a scheme to implement this gate in neutral atoms at the end of this work.
We envision a zoned architecture in which the gates are implemented in separate physical regions of the quantum processor, which has been experimentally realized for qubit-only neutral atom quantum computing in Ref. [38]; we specify the zones in which each of the gates is performed in Fig. 1b, and introduce our circuit notation.
To define our codes, it is convenient to introduce Majorana fermion operators, which are the real and imaginary parts $\gamma_{i}$, $\tilde{\gamma}_{i}$ of complex fermionic operators, $p_{i}^{\dagger}=\frac{1}{2}(\gamma_{i}+i\tilde{\gamma}_{i})$, fulfilling $\{\gamma_{i},\gamma_{j}\}=2\delta_{ij}=\{\tilde{\gamma}_{i},\tilde{\gamma}_{j}\}$ and $\{\gamma_{i},\tilde{\gamma}_{j}\}=0$. We construct codes in which we encode two logical Majorana fermions $\gamma^{\mathrm{L}}$ and $\tilde{\gamma}^{\mathrm{L}}$. These are then combined into a logical complex fermion $c^{\dagger}=\frac{1}{2}(\gamma^{\mathrm{L}}+i\tilde{\gamma}^{\mathrm{L}})$. In order for logical operators between different codeblocks to fulfill Majorana fermion anti-commutation relations, they each must be a product of an odd number of $\gamma_{i}$ and $\tilde{\gamma}_{i}$, i.e. they must have odd operator weight.
Repetition code for phase errors
We start with a repetition code which only corrects for phase errors $E=n_{i}$. This is the dominant error source in neutral atoms [36]. As an introductory example, consider the two-site code
|$\ket{0}_{\mathrm{L}}$ $=\frac{1}{\sqrt{2}}\left(\ket{00}-\ket{11}\right),$ (5)
|$\ket{1}_{\mathrm{L}}$ $=\frac{1}{\sqrt{2}}\left(\ket{10}+\ket{01}\right),$ (6)
where |$\ket{11}=p_{2}^{\dagger}p_{1}^{\dagger}\ket{00}$ and $p_{1}^{\dagger}\ket{00}=\ket{01}$ and |$\ket{0}$, |$\ket{1}$ denotes the absence, presence of a fermion, respectively. The above code can detect a single phase error, but does not detect loss errors $E=p_{i}$. The code has two logical Majorana fermion operators $\gamma^{\mathrm{L}}=\gamma_{1}$ and $\tilde{\gamma}^{\mathrm{L}}=\tilde{\gamma}_{2}$, from which we form a logical complex fermion operator by $c^{\dagger}=\frac{1}{2}\left(\gamma_{1}+i\tilde{\gamma}_{2}\right)$. This operator acts as $c^{\dagger}\ket{0}_{\mathrm{L}}=\ket{1}_{\mathrm{L}}$, $c^{\dagger}\ket{1}_{\mathrm{L}}=0$. We therefore interpret |$\ket{0}_{\mathrm{L}}$ and |$\ket{1}_{\mathrm{L}}$ as logical fermionic Fock states. Two codeblocks $a$ and $b$ each encode such an operator and they fulfill $\{c^{\dagger}_{a},c_{b}^{\dagger}\}=0$. Hence, $c^{\dagger}$ is indeed a logical fermion operator. This simple two-site code is also a stabilizer code, see Fig. 2a: its codestates are defined by the operator $i\tilde{\gamma}_{1}\gamma_{2}$ which acts as the unity operator on both states, $i\tilde{\gamma}_{1}\gamma_{2}\ket{0}_{\mathrm{L}}=\ket{0}_{\mathrm{L}}$ and $i\tilde{\gamma}_{1}\gamma_{2}\ket{1}_{\mathrm{L}}=\ket{1}_{\mathrm{L}}$. This enables generalization to detecting $N-1$ phase errors by using $N$ complex physical sites and the stabilizers $i\tilde{\gamma}_{i}\gamma_{i+1}$ for $1<i<N$. For $N\geq 3$, we can also correct phase errors as we show below. This code can also be used as a purely error-detecting code for all errors except atom-loss errors on the edges. In fact, this is the code defined by the ground space of the Kitaev chain, and the two logical Majorana operators $\gamma_{1}$ and $\tilde{\gamma}_{N}$ are its well-known edge states used as qubits in Majorana nanowires [37]. However, by contrast to nanowires, we do not implement the Kitaev Hamiltonian in our system. Hence, we cannot rely on the gap between the ground space of the Hamiltonian and the excited states in order to suppress errors and keep the computation in the ground space of the Hamiltonian. Instead, we require active error correction.
Figure 2: Logical operations for a repetition code protecting against phase errors. a, Illustration of splitting physical fermions for $N=3$ into virtual Majorana fermions. Stabilizers are given by $i\tilde{\gamma}_{1}\gamma_{2}$, $i\tilde{\gamma}_{2}\gamma_{3}$, and the logical fermionic creation operator is given by $c^{\dagger}=\frac{1}{2}(\gamma_{1}+i\tilde{\gamma}_{3})$. b, Syndrome measurement and correction. The two stabilizers are measured by using a Hadamard test on qubits. Conditioned on the measurement outcome, gates are performed to correct for errors. c, Encoding circuit, translated from Ref. [39], realized as a chain of braidings. Note that $(\mathrm{S}^{\mathrm{f}})^{\dagger}$ can be interpreted as intra-site braiding. d, Logical gates. $\sqrt{\mathrm{BRAID}}_{31}=\exp\left(-\frac{\pi}{8}\tilde{\gamma}_{3}\gamma_{1}\right)$, which only acts on the two edge atoms. See Methods for a derivation of these circuits.
For implementing active error correction, we need to measure where errors happened and correct for them, which we show how to do in Fig. 2b. We use our qubit-fermion gate to map a measurement of the fermion density $n$ to a measurement of the phase of the qubit, which we transform to a measurement of the population of the qubit using Hadamard gates. To measure the stabilizer instead of $n$ on the fermions we conjugate the qubit-fermion gate with two braiding gates as well as phase gates. An ancilla qubit is used for each of the stabilizers. Dependent on the measurement outcome, $\mathrm{Z}^{\mathrm{f}}$  gates are performed on the fermions to correct for the errors. This scheme shows the first advantage of using qubits: destructive qubit readout is faster than a non-destructive $n$ measurement [40]. To prepare a logical codestate |$\ket{0}_{\mathrm{L}}$, we use a chain of braiding gates—the $\mathrm{S}^{\dagger}$ gates effectively braid the two Majoranas on the same physical site, Fig. 2c.
Next, we discuss how to implement logical gates. We show the circuits in Fig. 2d, see Methods for a proof. The basic building block for the logical $\mathrm{CZ}^{\mathrm{f}}$ is the physical $\mathrm{CZ}^{\mathrm{f}}$. Similar to the stabilizer measurement, conjugation with braiding gates “spreads” the entanglement induced by the physical $\mathrm{CZ}^{\mathrm{f}}$ to the correct site, with $\mathrm{S}^{\mathrm{f}}$ gates selecting whether $\tilde{\gamma}$ or $\gamma$ is acted upon. The braiding gate is simple as it amounts to a physical braiding gate between $\gamma_{1}$ of one block and $\tilde{\gamma}_{N}$ of the other. The $\mathrm{S}^{\mathrm{f}}$ gate amounts to a braiding gate between the first and $N$th site of a single codeblock. Similarly, the $\mathrm{T}^{\mathrm{f}}$ gate is implemented as a braiding operation with half the duration of the $\mathrm{S}^{\mathrm{f}}$ gate. Finally, the qubit-fermion $\mathrm{CZ}^{\mathrm{qf}}$  is implemented similarly to the $\mathrm{CZ}^{\mathrm{f}}$ by conjugating the physical $\mathrm{CZ}^{\mathrm{qf}}$  with braid and phase gates.
The utility of this code is based on the fact that, in neutral atoms, phase errors are orders of magnitude more frequent than atom-loss errors. However, because braiding gates convert a single phase error into two correlated atom-loss errors, which cannot be detected, the depth of the circuits which can be performed in practice is still limited by the phase coherence time. Hence, this code is not fault-tolerant even in the absence of atom-loss errors. The fact that logical gates only act on the edge fermions on the other hand reduces the depth of the circuits, as the gate overhead for implementing logical gates is independent of $N$. It might be beneficial to reduce gate depth further by replacing the $\mathrm{CZ}^{\mathrm{f}}$gates and BRAIDs in Fig. 2d with their arbitrary-angle versions, which turns our gates to arbitrary-angle logical gates. However, for large-scale computations, a truly fault-tolerant code is necessary.
Fermion-qubit fault tolerance
We now show how to construct an error-correcting code correcting for both phase errors $E=n$ and atom-loss errors $E=p$ and show how to perform fault-tolerant gadgets. In the context of encoding qubits in Majorana nanowire platforms, Bravyi, Terhal, and Leemhuis [35] noted that a particular class of qubit error correcting codes, so-called weakly self-dual Calderbank-Shor-Steane (CSS) codes, can be converted into a Majorana stabilizer code by replacing qubit operators in the stabilizers by Majorana fermions. Motivated by the prospect of high error-resilience for codes with odd-weight logical operators, they noted that for codes with an odd number of physical Majorana fermions, a single odd-weight logical operator can be encoded. This operator is given by the total parity, i.e. the product of all Majorana operators. Our key observation is that, because there is an odd number of physical Majorana fermions, this logical operator is in fact a logical Majorana fermion operator. Even more importantly, while in Majorana nanowire platforms, a complex physical fermion would have to be encoded in two logical Majoranas which are encoded in spatially distinct physical Majoranas, our framework using complex physical fermions enables the encoding of two such Majorana codes, and therefore one complex logical fermion, in the same set of physical fermions. It turns out that a general prescription to constructing complex fermion codes from weakly self-dual CSS codes is to replace the $Z$ operators in the qubit code stabilizers by $\gamma$ operators and the qubit $X$ operators by $\tilde{\gamma}$ operators, see Fig. 3a. The logical Majorana operators are then given by the two parities $\gamma^{\mathrm{L}}=\prod_{i=1}^{N}\gamma_{i}$ and $\tilde{\gamma}^{\mathrm{L}}=\prod_{i=1}^{N}\tilde{\gamma}_{i}$, from which we construct a single logical fermion operator $c^{\dagger}=\frac{1}{2}\left(\gamma^{\mathrm{L}}+i\tilde{\gamma}^{\mathrm{L}}\right)$. As an example, we consider the family of triangular honeycomb Majorana color codes, which encodes a single logical fermion with odd distance $d$ into $N=(3d^{2}+1)/4$ physical fermions. In particular, $d=3$ is the fermionic version of the Steane code [41], Fig. 3a. Such codes can correct $(d-1)/2$ loss or phase errors.
Figure 3: Fermion-qubit fault-tolerant quantum computation. a, Construction of the fermionic Steane code, the smallest member of the family of color codes we consider. Stabilizers are obtained by replacing products of $X$ ($Z$) operators in the qubit code with products of $\gamma$ ($\tilde{\gamma}$), respectively. Corners on areas with the same color are part of a stabilizer as shown for the bottom left area. b, Mid-circuit measurement of $-\gamma_{1}\gamma_{2}\gamma_{3}\gamma_{4}$ (counting sites from bottom to top). By inverting the $\mathrm{BRAID}$s (i.e. $\mathrm{BRAID}_{12}\rightarrow\mathrm{BRAID}_{21}$), we measure $-\tilde{\gamma}_{1}\tilde{\gamma}_{2}\tilde{\gamma}_{3}\tilde{\gamma}_{4}$, the other stabilizer. c, Neutral-atom implementation of transversal $\mathrm{CZ}^{\mathrm{qf}}$  gates, performed by interleaving a logical qubit and fermion and applying a global Rydberg laser. This implements the circuit shown in subfigure e. d, Transversal $\mathrm{CZ}^{\mathrm{f}}$. This is implemented by the logical qubit in subfigure c with a logical fermion. e, Transversal qubit-fermion $\mathrm{CZ}^{\mathrm{qf}}$. Transversal f, $\mathrm{BRAID}$  and g, $\mathrm{S}^{\mathrm{f}}$  gates are similarly implemented by parallel operations. h, $\mathrm{T}^{\mathrm{f}}$gate. Implemented by using a T gate applied on an ancilla qubit, which is then mapped onto the fermion. All the operations shown here are straightforwardly generalized to codes with larger (odd) $d$, see Methods.
Of central utility is the fact that we can use qubits for controlled operations. One example is the measurement of stabilizers using the Hadamard test, as shown in Fig. 3b. From these stabilizers, we can reconstruct whether a $\tilde{\gamma}$ or $\gamma$ error (or both) happened on a fermion. Such single-Majorana errors are corrected by coupling an ancilla site $a$ prepared in |$\ket{0}$ to the site with the error, acting with $(\mathrm{BRAID}_{ai})^{2}=\gamma_{i}\tilde{\gamma}_{a}$ and then measuring the ancilla. The measurement circuit in Fig. 3b is made fault-tolerant by for example using the Shor scheme, i.e. replacing the single ancillary qubit with a set of qubits in a cat state and repeatedly measuring the stabilizers. This stabilizer measurement is also used for fault-tolerant preparation of the logical codestates from a trivial starting state. Note that the correction of stabilizers is optional for both state preparation and error correction—instead of enforcing stabilizers to be +1, their value can instead be tracked [42]. Moreover, we expect that error correction only needs to be applied after each gate, and not during movement, as demonstrated in neutral-atom qubit quantum computers [38].
Magically, this fermion Steane code yields highly intuitive and easy-to-implement fault-tolerant logical gates. Transversal logical gates are implemented, similar to qubit Steane codes, by applying the corresponding physical gate in parallel on all atoms, c.f. Fig. 3c-g, see Methods. This set of gates turns out to be the Majorana Clifford gates $\mathrm{BRAID}$, $\mathrm{CZ}^{\mathrm{f}}$, $\mathrm{S}^{\mathrm{f}}$, which in analogy to qubit Clifford gates map a single Majorana operator string to another single Majorana string. These transversal Clifford gates are inherently fault-tolerant as an error on one physical codeblock can only spread to one physical fermion in the other codeblock.
One of the key tools enabled by our approach is the interfacing of our fermionic Steane code with a qubit Steane code (with matching $N$) using the qubit-fermion $\mathrm{CZ}^{\mathrm{qf}}$ gate: by applying physical $\mathrm{CZ}^{\mathrm{qf}}$ in parallel (see Fig. 3e), we realize a logical, transversal $\mathrm{CZ}^{\mathrm{qf}}$. This enables fault-tolerant qubit-fermion quantum computation. In particular, we will discuss in the next section that this has immense utility for implementing modern quantum algorithms for simulating fermions.
This gate set of $\mathrm{BRAID}$, $\mathrm{CZ}^{\mathrm{qf}}$, $\mathrm{CZ}^{\mathrm{f}}$, $\mathrm{S}^{\mathrm{f}}$, together with $\mathrm{H}$ and $\mathrm{S}$ on qubits, forms the “qubit-fermion Clifford group” as it maps a Majorana-qubit string to a single other Majorana-qubit string. This gate set needs to be combined with a $\mathrm{T}^{\mathrm{f}}$gate or a qubit T gate to enable universal qubit-fermion computation. While Majorana fermion schemes for T-gate synthesis have been proposed [43], they are more involved than qubit schemes due to the requirement to create two-fermion magic states. We instead propose using a qubit $\mathrm{T}$ gate applied on a qubit Steane code and transferring it to the fermions using our transversal logical $\mathrm{CZ}^{\mathrm{f}}$gate, Fig. 3h. This therefore enables a significant simplification compared to Majorana $\mathrm{T}$-gate synthesis.
Logical qubit-fermion quantum simulation
Qubits Fermions
fSWAP [44] pFFFT+pINT FFFT+pINT
Depth $\mathcal{O}(N)$ $\mathcal{O}(\log{N})$ $\mathcal{O}(\log{N})$
Ancillas 0 $\mathcal{O}(N^{2})$ + $\mathcal{O}(N\log{N})$ 0 + $\mathcal{O}(N\log{N})$
Volume $\mathcal{O}(N^{2})$ $\tilde{\mathcal{O}}(N^{2})$ $\tilde{\mathcal{O}}(N)$
Figure 4: Fermion-qubit logical quantum simulation primitives and comparison to the qubit-only approach. a, Qubit-controlled arbitrary-angle interaction gate. The box with label “$\theta/4$” is a qubit gate $\exp(-i(\theta^{\prime}/2)X)$ with $\theta^{\prime}=\theta/4$, synthesised from discrete gates using the Solovay-Kitaev algorithm. See Fig. 1 for our convention of the labeling of the braiding gate. d, Qubit-controlled arbitrary-angle braiding gate. The gray box is the qubit-controlled braiding gate $\exp(-(\pi/4)Z\tilde{\gamma}_{i}\gamma_{j})$. c, FFFT circuit implemented using logical fermions. Crossing lines indicate swapping, implemented by movement of the atoms. d, Corresponding FFFT circuit implemented using logical qubits, c.f. Ref. [45]. e, Decomposition of the n-mode FFFT into an arbitrary-angle phase gate $\exp(-i\theta n)$ with $\theta=\theta_{k}^{n}=2\pi k/n$ and a two-mode FFFT. f, Compilation of the two-mode FFFT into fermionic fault-tolerant gates. g, Scaling of depth, number of ancillas and circuit volume (defined as the product of depth and the number of ancillas) with the number of sites $N$ for a single Trotter step of the plane-wave basis Hamiltonian in Eq. (7). We compare the lowest-depth previous method [12], our qubit-only and our fermion-qubit algorithms. “pINT” refers to a parallelized interactions scheme using the quantum FFT [46], “pFFFT” to our version of the FFFT which using the parallelization of CZ circuits introduced in Ref. [47].
A prime area of application for fermionic quantum computing is simulating electrons in molecules and materials. In second quantized approaches, the continuous-space wavefunctions of electrons are cast into a discrete lattice model through a choice of basis, which leads to a Hamiltonian coupling fermionic operators as $H_{\mathrm{qchem}}=\sum_{ij}J_{ij}c^{\dagger}_{i}c_{j}+\sum_{ijkl}V_{ijkl}c^{\dagger}_{i}c^{\dagger}_{j}c_{k}c_{l}$, with $J_{ij}$ and $V_{ijkl}$ constants that depend on the choice of basis. This generally leads to a number of Hamiltonian terms scaling as $\sim N^{4}$ with the basis size $N$. For a Trotterized simulation, this implies the same scaling for the number of gates per Trotter step. However, the plane-wave-dual basis [12], which is particularly suitable for simulating crystalline materials, enables writing the chemistry Hamiltonian as
$H$ $=\mathrm{FFFT}^{\dagger}\sum_{i}\epsilon_{i}n_{i}\mathrm{FFFT}$
$\quad+\sum_{i<j}U_{ij}\left(n_{i}-\frac{1}{2}\right)\left(n_{j}-\frac{1}{2}\right),$ (7)
where $\epsilon_{i}$, $\mu_{i}$, and $U_{ij}$ are classically calculable coefficients and $\mathrm{FFFT}$ denotes the fermionic fast Fourier transform  [45]. This formulation reduces the scaling of the number of terms to $\sim N^{2}$. Efficient algorithms for dynamics, as well as for ground and finite-temperature state preparation, in addition often require controlling this Hamiltonian on a qubit [10, 11].
Our next goal is therefore to implement qubit-controlled arbitrary-angle evolution under each of the terms in Eq. (7) individually. Time evolution under the Hamiltonian $\epsilon_{i}n_{i}$ until time $t$ is implemented by replacing the T gate in Fig. 3h with an approximate arbitrary-angle rotation of angle $\epsilon_{i}t$, where $t$ corresponds to time, synthesized from discrete gates with e.g. the Solovay-Kitaev algorithm. The interaction gate $\exp\left(-iZ\frac{\theta}{2}(n_{i}-\frac{1}{2})(n_{j}-\frac{1}{2})\right)$ is implemented by the circuit shown in Fig. 4a, which maps the arbitrary-angle single-qubit gate onto the fermions using $\mathrm{CZ}^{\mathrm{qf}}$  gates. A related gadget shown in Fig. 3b realizes an arbitrary-angle qubit-controlled braiding gate $\exp(-\frac{\theta}{2}Z\tilde{\gamma}_{i}\gamma_{j})$, where again $\mathrm{CZ}^{\mathrm{qf}}$  gates map the phase gate from the qubit to a fermion. The $\mathrm{BRAID}$s then “spread” the gate onto the second fermion. Two such gadgets are combined to yield the qubit-controlled hopping gate via $\exp\left(iJ_{ij}\Delta tZ\left(c_{i}^{\dagger}c_{j}+h.c.\right)\right)=\exp(-\frac{J\Delta t}{2}Z\tilde{\gamma}_{j}\gamma_{i})\exp(-\frac{J\Delta t}{2}Z\tilde{\gamma}_{i}\gamma_{j})$. These arbitrary-angle gates also enable time evolution under the more general Hamiltonian $H_{\mathrm{qchem}}$ by using the compilations in Ref. [36]. Finally, we need to implement the FFFT, which may be reduced to the 1D FFFT circuit shown in Fig. 4c, c.f. Ref. [45]. In analogy to the classical fast Fourier transform, the FFFT uses a divide-and-conquer reduction to two-site Fourier transforms $F^{k}_{N}$ (Fig. 4e, f) whose phases depend on the number of sites $N$ and an index $k$. Despite the fact that it is a 1D transformation, it requires non-local connectivity, which is implemented in neutral atoms by moving the tweezers [48].
Advantages over qubit quantum simulation
We now show that our approach yields a large advantage over currently known qubit algorithms for simulating Trotterized time evolution under the Hamiltonian of crystalline materials.
[…]
By contrast, in our fermionic platform, the fact that fSWAP gates are implemented by atom rearrangements (Fig. 4c) means that a depth of $\mathcal{O}(\log N)$ can be implemented without ancillas, yielding $\mathcal{O}(N)$ space and $\tilde{\mathcal{O}}(N)$ volume complexity. We note that the $N$-scaling of movement time does not prevent the exponential depth reduction from leading to an exponential time saving. This is because the constant overhead of error correction dominates, see Methods. We achieve another depth improvement through our compilation of the constituent FFFT subroutines (see Fig. 4e, f), which only takes half the previously-known T-gate depth [13].
At first sight, the logarithmic depth of the kinetic term is not useful because the interaction term requires $\mathcal{O}(N)$ depth because of the parallelizability restriction imposed by the $\mathcal{O}(N^{2})$ terms. However, similarly to the qubit version of the FFFT, we can use ancillas to reduce the circuit depth by using the translational invariance of $U_{ij}$. This enables using a quantum circuit of the classical fast Fourier transformation [46, 49], see Methods. Because the Fourier transform of $n_{i}$ is a complex number, we encode it in a binary register with $\mathcal{O}(\log(N))$ ancillas. Therefore, this “parallelized interactions” scheme in total uses $\mathcal{O}(N\log(N))$ ancillas, for both qubit and fermion quantum computers. After this transform, the density interactions are diagonal and implementable in depth $\mathcal{O}(\mathrm{log}(N))$.
Taking both interaction and kinetic terms together, we thus achieve a log-depth Trotter step for materials. For qubit quantum computers, it uses $\mathcal{O}(N^{2})$ qubits, but for fermionic quantum computers only $\mathcal{O}(N\log(N))$ fermionic sites. Therefore, the circuit volume is a factor of $N/\log(N)$ lower for fermions.
The remaining question is how this log-depth Trotter step reflects on the end-to-end complexity of time-evolving up to time $t$. We use the Trotter scheme introduced in Ref. [50], which uses $\mathcal{O}(Nt)$ Trotter steps for fixed simulation accuracy and large, fixed Trotter order. Using our log-depth Trotter steps, we therefore find an overall depth complexity of $\tilde{\mathcal{O}}(Nt)$, which is essentially linear in both time and the number of sites. While we show that both qubit and fermion quantum computers can reach this scaling, only the fermion version requires a linear number of modes, while at least within our approach, qubits need a prohibitively expensive quadratic number of qubits using fSWAP networks.
For the more general $H_{\mathrm{qchem}}$, at first glance, no depth advantage appears possible, since current state-of-the-art schemes combining double factorization in the molecular basis with fSWAP networks achieve depth $\mathcal{O}(N^{2})$ without ancillas [51] and $\mathcal{O}(N)$ depth with ancillas [52] for a single Trotter step, independent of the molecule size. By contrast, fermionic quantum computers can reach depth $\mathcal{O}(N)$ for large molecules, see Methods, which to our knowledge is the first algorithm for time evolution under the chemistry Hamiltonian which achieves linear depth per Trotter step without using ancillas. For Hamiltonians which do not allow for a low-rank factorization, such as the sparse Sachdev-Ye-Kitaev model [53], we expect such depth speedups to become apparent already at small system sizes.
Within our fault-tolerant gate set, no advantage in the number of T gates is possible if the T-gate count is the only parameter which is optimised for. This can be directly seen by mapping the fermionic gates to qubit gates using the Jordan-Wigner transform and noting that only the $\mathrm{T}^{\mathrm{f}}$ gate requires T gates (Methods). We note that steady progress in T-gate synthesis means they do not completely dominate the simulation cost over Clifford gates [34]. Moreover, our depth reductions directly translate into a large space-time volume saving compared to qubit platforms. Therefore, we expect advantages in terms of the total runtime of a fermionic simulation.
Figure 5: Implementation of the braiding gate in neutral atoms. a, Illustration of the sequence of operations to implement an effective braiding gate. Braiding is a product of pairing and the tunneling gate $\sqrt{\mathrm{iSWAP}}^{\mathrm{f}}=U_{t\downarrow}(\pi/4)$. $U_{\mathrm{t}\downarrow}(\theta)$ and $U_{\uparrow\downarrow}(\theta)$ act only on fermions, while the beamsplitter acts only on molecules, and $U_{\mathrm{diss}}$ is a fermion-molecule gate. The pairing gate dissociates a bosonic molecule into two fermions on two separate sites. The pulse duration of the dissociation interaction depends on the average number $M=\braket{b^{\dagger}b}$ of molecules in the tweezer. b, The molecules are formed from the same atomic species as the fermionic atoms used. c, Ramsey sequence to probe the effective coherence of a “Bell-pair” (|$(\ket{00}+i\ket{11})/\sqrt{2}$ created by the effective pairing gate. The phase gate is defined as $\exp(-i\theta n)$. The beamsplitter gate, defined above Eq. (9), has phase $\theta=\pi/4$. d, Ramsey fringes as a function of initial molecule Fock state number $M$. e, Simulated pairing gate infidelity as a function of molecule Fock state number $M$. See Fig. S1f for the circuit used to determine the infidelity.
Implementation in neutral atoms
Finally, we describe how to implement our framework in neutral atoms, which requires the implementation of the universal qubit-fermion gateset shown in Fig. 1. $\mathrm{CZ}^{\mathrm{qf}}$ and $\mathrm{CZ}^{\mathrm{f}}$ are implemented using a high-lying Rydberg state of the atom, similarly to the routinely implemented qubit-qubit $\mathrm{CZ}^{\mathrm{q}}$ [54, 36]. $\sqrt{\mathrm{iSWAP}}^{\mathrm{f}}$ is implemented by bringing two tweezers close to each other, inducing tunneling of atoms between the sites [36] or by placing the atoms into an optical lattice [55] and controllably performing tunneling by using superlattices [56, 57]. $\mathrm{T^{\mathrm{f}}}$, $\mathrm{S^{\mathrm{f}}}$, $\mathrm{Z^{\mathrm{f}}}$ are implemented by detuning the trapping frequency of the target tweezer or optical lattice site. What is left to show is how to implement $\mathrm{BRAID}$. We propose its realisation as a product $\mathrm{BRAID}_{ij}=\mathrm{PAIR^{\prime}}_{ij}\mathrm{\sqrt{iSWAP}}^{\mathrm{f}}_{ij}$, where $\mathrm{PAIR^{\prime}}_{ij}$ is the pairing gate
$\mathrm{PAIR^{\prime}}_{ij}=\exp\left(i\frac{\pi}{4}\left(p_{i}^{\dagger}p_{j}^{\dagger}+h.c.\right)\right).$ (8)
To implement pairing, we propose the pulse-sequence in Fig. 5a. It employs photodissociation of homonuclear diatomic bosonic molecules consisting of the same atomic species as the fermions [58, 59], see Fig. 5b. Photodissociation realizes the boson-fermion gate $U_{\mathrm{diss}}=\exp\left(-i\frac{\pi}{4\sqrt{M}}\left(b_{i}p^{\dagger}_{i\uparrow}p^{\dagger}_{i\downarrow}+b^{\dagger}_{i}p_{i\downarrow}p_{i\uparrow}\right)\right)$, where $b$ is the bosonic annihilation operator of molecules (we assume there are $M$ of them) residing in the ground state of the tweezer and obeying commutation relations $[b_{i},b_{j}^{\dagger}]=\delta_{ij}$. $\uparrow$ and $\downarrow$ label two internal states of the atom, where so far we assumed that all atoms reside in $\downarrow$, i.e. $p_{i}=p_{i\downarrow}$. In order to convert this same-site gate into a non-local gate, we conjugate it with a fermion spin-selective tunneling $U_{\mathrm{t}\downarrow}(\theta)=\exp\left(-i\theta\left(p^{\dagger}_{i\uparrow}p_{j\uparrow}+h.c.\right)\right)$ and a fermionic spin-flip $U_{\uparrow\downarrow}(\theta)=\exp\left(-i\theta\left(p^{\dagger}_{i\downarrow}p_{i\uparrow}+h.c.\right)\right)$. To suppress interactions between the molecules and the fermions, the molecules reside in a separate tweezer when not in use for gates and are tunneled into the target tweezer by moving the two tweezers close to each other, implementing a beam-splitter operation $\mathrm{BS}(\theta)=\exp\left(\theta\left(b_{i}^{\dagger}b_{j}-h.c.\right)\right)$. Altogether, the sequence in Fig. 5a implements
$\mathrm{PAIR}_{ij}=\exp\left(i\frac{\pi}{4\sqrt{M}}\left(b_{k}p_{i}^{\dagger}p_{j}^{\dagger}+h.c.\right)\right).$ (9)
We now discuss how we use Eq. (9) to effectively implement Eq. (8). If the molecules were in a coherent state with large amplitude $\sqrt{M}$, we could replace $b\rightarrow\braket{b}=\sqrt{M}$ and the two gates would be equal. However, in the case of atoms trapped in dipole traps, Fock states or mixtures thereof are more naturally prepared than coherent states. To overcome this conceptual challenge, we in the following use the fact that large-$M$ Fock states cannot be distinguished from coherent states by measurements [60].
To show that we can indeed effectively implement Eq. (8) employing a large-$M$ Fock state of molecules, consider the thought experiment shown in Fig. 5c. The first pairing gate is supposed to create a “fermionic Bell-pair” (|$(\ket{00}+i\ket{11})/\sqrt{2}$. To probe whether there is coherence, a phase gate with phase $\theta$ is applied and a second $\mathrm{PAIR}$ gate rotates the measurement basis into the effective Bell basis. This realizes a Ramsey-type experiment. Simulating this experiment numerically, we find Ramsey fringes with a contrast that approaches unity as $M\rightarrow\infty$. This shows that, after the first pairing gate, there was indeed effective coherence between states |$\ket{00}$ and |$\ket{11}$. An intuitive way to understand this observation is that the molecule number distribution in each mode after the beamsplitter is approximately Poissonian—just like in a coherent state. The entanglement between both modes then leads to an effective phase coherence between the modes. This effective coherence is then revealed by the second pulse, which disentangles the fermions from the molecules. We say “effective” coherence because the reduced density matrix of the fermions is always diagonal; indeed, by measuring the total number of molecules, one projects the fermions onto |$\ket{00}$ or |$\ket{11}$. Therefore, it is important that the molecules are not measured while the fermions are entangled with them. In Fig. S1, we additionally perform state tomography numerically, showing that the state after the first pairing gate is indeed effectively a Bell pair.
We probe the gate fidelity of our effective pairing gate in Fig. 5e using Choi state tomography, which we generalize to fermionic gates, see circuit shown in Fig. S1. We find that only $100$ molecules suffice to create a gate with $>99.8\,\%$ fidelity. In the Methods, we also show that even a mixed state with a Poissonian number distribution can be used instead of |$\ket{M}$. Non-trivially, Fig. 5c and Fig. 5d show that our scheme works even if the second pairing gate is not performed on the same molecule site, which enables full parallelizability of pairing operations. For that to be the case, the molecule tweezers need to be effectively coherent with respect to each other, just like lasers manipulating different qubits within a quantum computer need to be phase-locked. The cross-coherence of molecule tweezers is guaranteed by creating them from one common Fock state with a beamsplitter as done in Fig. 5c. To parallelize this further, many molecule tweezers can be created by repeatedly applying beamsplitters starting from one large Fock state, for example by tunneling from a single large dipole trap. While our proposal is challenging to implement, coherent dissociation of Feshbach molecules is routinely performed in cold-atom experiments [58].
Outlook
Our work opens the field of fermion-qubit error correction and fault-tolerant algorithms. Of prime importance is the error threshold of our codes for the strongly biased errors present in experiments. Moreover, fermionic versions of quantum low-density-parity check codes may be developed to enable a finite rate of logical fermions per codeblock. Due to our reliance on transversal gates, algorithmic fault tolerance can be explored within our approach [34].
We have so far mainly used the presence of qubits as a helper to implement logical operations. In high-energy physics, the possibility to directly simulate qubit-fermion models enables simulation of fermions coupled to large-spin degrees of freedom in lattice gauge theories [36, 61]. For this purpose, qudits could be directly implemented with large-spin atomic species [62, 63], rotational levels of molecules [64], or an oscillator mode [65]. Relatedly, the depth-efficiency of our approach might enable simulation of quantum gravity models [53].
Boson-fermion models could be studied by combining our schemes with oscillator-qubit quantum simulation [66, 61] using either motional modes of neutral atoms in tweezers [67, 61] or bosonic atoms. Suitable boson-to-boson error correction schemes would need to be found [68, 69]. This way, the longstanding problems of the influence of phonon-electron interactions on thermal, structural, and vibrational material properties [70] as well as the interplay of molecular vibrations with the electronic wavefunction in chemistry [71] could be investigated using potentially vastly shallower circuits than with qubit-only platforms.
Our architecture might offer an advantageous approach not only for simulating fermions but also for quantum computations that are natively qubit-only. We encoded logical fermion operators in odd-weight fermion operators—instead, logical qubits can be encoded in even-weight logical operators, for example Majorana surface codes which have been shown to have high thresholds [72]. One advantage of this approach is that lost atoms are replenished by coupling to the molecule bath, offering an alternative to continuous atom reloading [55], which is one of the challenging aspects of neutral-atom quantum computing.
Finally, motivated by our depth improvement of the qubit algorithm for materials, development of fermion-hardware-inspired qubit algorithms might lead to further improvements in qubit algorithms for fermionic simulation.
Acknowledgments
We acknowledge discussions with Ignacio Cirac, Daniel Gonzalez-Cuadra, Christopher Kang, Paul Julienne, Robert Ott, Shraddha Singh, Maryam Mudassar, Nathan Wiebe, Torsten Zache, Johannes Zeiher, and Hengyun Zhou.
Funding. This material is based upon work supported in part by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. Additional support is acknowledged from the following agencies. This work was supported in part by the Department of Energy (DOE), Office of Science, Office of Nuclear Physics, via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science (award no. DE-SC0023710) and the National Science Foundation’s Quantum Leap Challenge Institute in Robust Quantum Simulation (award no. OMA-2120757). A.V.G. was also supported in part by the DoE ASCR Accelerated Research in Quantum Computing program (award No. DE-SC0020312), AFOSR MURI, DoE ASCR Quantum Testbed Pathfinder program (awards No. DE-SC0019040 and No. DE-SC0024220), NSF STAQ program, and DARPA SAVaNT ADVENT. E.C. was also supported in part by ARO MURI (award No. SCON-00005095), and DoE (BNL contract No. 433702). Support is acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. C2QA participated in this research.
Competing interests. All authors declare no competing interests.